A sign (mass 2100 kg k g ) hangs from the end of a vertical steel girder with a cross-sectional area of 0.015 m2 m 2 . Ignore the mass of the girder itself. The elastic modulus for steel is 2.0×1011N/m2 2.0 × 10 11 N / m 2 .

A. What is the stress within the girder?

B. What is the strain on the girder?

C. If the girder is 7.20 mm long, how much is it lengthened?
Answer #1

Given: m= 2100 kg , g=9.8m/s² A=0.015 m² Y = 2x10N/ms 6 stress, mg (A) 2100x9.8 S- A Oools = 1.372 x 10° N/m? S= 1.372x/6N/